Phineas is pushing a 10.62-kg box of tools with a force of 45 N directed at an angle of 81° downward from the horizontal. How fast is the cart going after the acceleration from rest if their backyard is 31 cm away from the garage?


1
Expert's answer
2021-03-02T18:06:49-0500
"v^2=2as\\v^2=2s\frac{F\cos{45}}{m}\\v^2=2(0.31)\frac{45\cos{81}}{10.62}\\v=0.64\frac{m}{s}"